a purse contains $ 6.50 in quarters and dimes.there are 5 more quarters than dimes.find the number of quarters and dimes
Answer by stanbon(75887)

a purse contains $ 6.50 in quarters and dimes.there are 5 more quarters than dimes.find the number of quarters and dimes
--------
Quantity: q = d + 5
Value: 10d + 25q = 650 cents
------------------------------------
Modify:
2d - 2q = -10
2d + 5q = 130
------------------
Subtract and solve for "q":
7q = 140
q = 20 (# of quarters)
------
Solve for "d":
d = q - 5
d = 20-5
d = 15 (# of dimes)
----------------------------
Cheers,
Stan H.